A projectile is fired horizontally at a speed of 8 m/s from an 80-m-high cliff. Find (a) The velocity just before the projectile hits the ground.


1
Expert's answer
2021-11-15T17:54:57-0500

The velocity is


"v=\sqrt{v_x^2+v_y^2},\\\space\\
v=\sqrt{v_x^2+(\sqrt{2gh})^2}=40\text{ m/s},\\\space\\
\theta=\arctan\frac{\sqrt{2gh}}{v_x}=89.6°"